OpenStudy (anonymous):

]Choose the equation of the line passing through the point (4, -2) and perpendicular to y = one thirdx - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (anonymous):

it wants to be peripendicular to the line so its slope should be -3 y-(-2)=-3(x-4) y=-3x+10